# Modulate and Reconstruct: Learning Hyperspectral Imaging from Misaligned Smartphone Views

### v3

@@ 5 Method
The task of MI-HSR involves predicting a hyperspectral image aligned to a target viewpoint (the Main camera), given RGB images from multiple spatially offset sensors:
$\mathbf{I}_{\mathrm{Main}},\mathbf{I}_{\mathrm{Tele}},\mathbf{I}_{\mathrm{Wide}}\in\mathbb{R}^{h\times w\times 3}.$
@@
This setup introduces two key challenges: (i) the input views are misaligned due to differing camera geometries, and (ii) the available hyperspectral supervision $\mathbf{I}_{\mathrm{HS}}\in\mathbb{R}^{h\times w\times n}$ corresponds to a reference sensor not aligned with any RGB input. Our approach addresses both issues, one through pre-processing, the other through architectural design.
5.1 Supervision Warping via Learned Optical Flow
@@
To leverage the hyperspectral reference image for training, we align it to the Main RGB view using learned OF. Since spectral and RGB images differ in modality, we first compute a color projection $\mathbf{C}\in\mathbb{R}^{n\times 3}$ to transform $\mathbf{I}_{\mathrm{HS}}$ into an RGB approximation:
$\mathbf{C}:=\operatorname*{argmin}_{\mathbf{C}}\|\mathbf{I}_{\mathrm{HS}}\mathbf{C}-\mathbf{I}_{\mathrm{Main}}\|_{2}^{2}.$
Given $\mathbf{I}_{\mathrm{HS}}\mathbf{C}$ and $\mathbf{I}_{\mathrm{Main}}$, we estimate a dense correspondence field $\bm{\delta}\in\mathbb{R}^{h\times w\times 2}$ using a pre-trained OF model $\mathcal{F}$ (PWC-Net [41]):
$\bm{\delta}:=\mathcal{F}(\mathbf{I}_{\mathrm{HS}}\mathbf{C},\mathbf{I}_{\mathrm{Main}}).$
@@
This flow is used to warp the hyperspectral GT to the Main view:
$\mathbf{I}_{\mathrm{HS}}^{\mathrm{w}}:=\mathcal{W}(\mathbf{I}_{\mathrm{HS}},\bm{\delta}),\quad\mathbf{m}:=\lfloor\mathcal{W}(\mathbf{1},\bm{\delta})\rfloor,$
@@
where $\mathbf{m}$ is a binary mask indicating valid visible pixels. This enables aligned supervision for training and pointwise loss computation $\mathcal{L}_{1}(\mathbf{I}_{\mathrm{HS}}^{\mathrm{w}}\circ\mathbf{m},\hat{\mathbf{I}}_{\mathrm{HS}}\circ\mathbf{m})$
5.2 Implicit Cross-View Alignment in Network Design
Even with warped supervision, the three input views remain spatially misaligned. Direct calibration or flow-based alignment is possible but impractical in general-purpose settings. Instead, we encode alignment into the architecture itself, drawing inspiration from recent findings that transformer attention can perform implicit alignment across modest viewpoint shifts [42].
We adopt MSFN [29] as our base model, which consists of spatial and spectral transformer-based U-Nets. However, its original design applies spectral modeling before spatial, which is unsuitable for misaligned inputs as spectral attention assumes spatial coherence across input channels, which multi-camera setups violate.
5.3 Proposed Architecture: MI-MSFN
We propose a revised architecture, Multi-Image MSFN (MI-MSFN), specifically tailored to address input misalignment and heterogeneous spatial content. Here, we reverse the order of modules, applying spatial transformer blocks before spectral ones (Figure 6). This allows the network to first establish spatial correspondences across the Main, Tele, and Wide inputs thus effectively aligning features by attention before fusing spectral information. We also remove the skip paths to enforce that all feature propagation passes through alignment-aware components.
Figure 6: Overview of the proposed MI-MSFN architecture for multi-image hyperspectral reconstruction. It integrates implicit alignment across misaligned inputs via spatial-first attention, followed by spectral modeling. For detailed module definitions (Spectral FM, Spatial FM), refer to Sec. A.4
@@ 6.1 Experimental Setup
Dataset. All the experiments were conducted on our Doomer dataset as it is the only available one for MI-HSR task. Experiments on single-image HSR were only conducted for comparison to MI-HSR, thus featuring the same dataset.
Metrics. We adopt well-known Peak Signal-to-Noise Ratio (PSNR) and Spectral Angle Mapper (SAM). We also introduce Normalized Spectral Error (NSE), which serves as an alternative to MRAE as it does not penalize having few dark bands and reflects the integral nature of radiance:
@@
When comparing predictions to warped GT, some pixels were missing or invalid, so we masked out positions at $\mathbf{m}(x,y)=0$.
Implementation details. We split our dataset into train and test subset in proportion 4:1. In all the experiments we used Adam optimizer with learning rate of 0.0004 inherited MSFN (except for AWAN at Sec. 6.3, which showcased the necessity of polynomial scheduler with power of 1.5 starting at learning rate of 0.0001). We trained all the networks on random $64~\times~64$ patches for $10^{4}$ epochs. Every 30 epochs we ran evaluation loop; the final model is the one with the best mean absolute error on the aligned test set. Each training procedure was re-run 10 times with different random seeds, yielding 10 distinct quantitative results. We summarize them by reporting mean and standard deviation.
@@ 6.2 Results
We trained the proposed MI-MSFN network using two input configurations: single-image $(\mathbf{I}_{\mathrm{Main}})$ and multi-image $(\mathbf{I}_{\mathrm{Main}},\mathbf{I}_{\mathrm{Tele}},\mathbf{I}_{\mathrm{Wide}})$. To validate our supervision alignment (Sec. 5), following [35], we report metrics both on aligned and original GT in Tab. 2. Our multi-image acquisition system improves HSR by 3.56 dB PSNR, 38% SAM, 28% NSE. The lower standard deviation (1–3% of mean value) underlines the learning stability of the proposed MI-MSFN approach.
Setting Aligned GT Original GT
PSNR, dB ↑ SAM, ° ↓ NSE, % ↓ PSNR, dB ↑ SAM, ° ↓ NSE, % ↓
Single-camera 26.30 ± 0.68 6.11 ± 0.40 12.71 ± 0.46 25.05 ± 0.48 6.25 ± 0.42 14.14 ± 0.67
Multi-camera 29.86 ± 0.21 3.77 ± 0.09 9.14 ± 0.20 27.88 ± 0.15 3.91 ± 0.08 10.35 ± 0.16
Table 2: Evaluations of MI-MSFN in single- and multi-camera settings. Mean and standard deviation of each metric is computed across 10 re-runs. The use of auxiliary cameras allows highly better HSR.
Fig. 7 presents a qualitative comparison between single-camera and multi-camera configurations. We show patch-level comparisons at a selected wavelength as well as radiance profiles at specific points of interest. The multi-camera system consistently recovers finer surface details that the single-camera setup fails to capture. In the first row, a printed symbol on a book cover is not registered by the multi-camera system at $\lambda=700~\text{nm}$ which complies with GT while SI-HSR falsely reveals the symbol. In the second row, the single-camera system introduces spurious noise-like structures that do not correspond to any physical features. In the third row, the single-image setup fails to reconstruct radiometric intensity accurately, significantly misestimating the radiance of the paper sheet. We also showcase a failure case of our system in the fourth row. These results highlight the improved spatial and spectral fidelity enabled by multi-view fusion in the MI-HSR system.
@@
Figure 7: Qualitative comparison of MI-MSFN predictions under single-image and multi-image input settings. For each scene, a specific spectral band, region of interest, and point of interest are selected. Regions are enlarged and displayed side-by-side to highlight differences in spatial and spectral reconstruction quality. Note that the grayscale renderings correspond to individual spectral bands, as shown at the wavelength indicated on the left of each image.
@@ 6.3 Ablation Study
We conducted two ablation studies, beginning with replacement of HSR network. We compared MI-MSFN against several approaches: HSCNN+ [24], AWAN [22, 43], MST++ [8] (winners of NTIRE 2018 [21], 2020 [22], 2022 [23] respectively) and MSFN [29] as shown in Tab. 4. Since we evaluate them in our misaligned multi-image setting they were not designed for, we also help them by warping inputs onto the Main camera view using the same pre-trained OF as before. In particular, Tab. 4 depicts the superiority of MI-MSFN over MSFN, justifying the proposed architectural changes.
Method OF PSNR, dB ↑ SAM, ° ↓ NSE, % ↓
HSCNN+ [24] 26.28 ± 0.40 5.66 ± 0.13 12.20 ± 0.13
26.46 ± 0.64 5.76 ± 0.13 12.23 ± 0.17
27.01 ± 0.79 4.56 ± 0.08 10.80 ± 0.10
AWAN [43] 26.41 ± 1.30 4.49 ± 0.17 10.68 ± 0.20
MST++ [8] 29.01 ± 0.25 4.18 ± 0.15 9.86 ± 0.17
29.37 ± 0.31 4.11 ± 0.07 9.66 ± 0.11
29.18 ± 0.21 3.91 ± 0.13 9.77 ± 0.13
MSFN [29] 29.59 ± 0.21 3.93 ± 0.13 9.33 ± 0.23
MI-MSFN (ours) 29.86 ± 0.21 3.77 ± 0.09 9.14 ± 0.20
Table 3: Ablation study on HSR model showing MI-MSFN being the best choice for MI-HSR.
Active cameras
Main Tele Wide PSNR, dB ↑ SAM, ° ↓ NSE, % ↓
26.30 ± 0.68 6.11 ± 0.40 12.71 ± 0.46
26.34 ± 0.67 5.77 ± 0.31 12.41 ± 0.46
29.63 ± 0.28 3.97 ± 0.11 9.08 ± 0.33
29.86 ± 0.21 3.77 ± 0.09 9.14 ± 0.20
Table 4: Ablation study on cameras configuration. Each camera brings an improvement in SAM.
The second ablation study evaluates contribution of each auxiliary camera. It extends Tab. 2 by two more settings: $(\mathbf{I}_{\mathrm{Main}},\mathbf{I}_{\mathrm{Tele}})$ and $(\mathbf{I}_{\mathrm{Main}},\mathbf{I}_{\mathrm{Wide}})$. Results are given in Tab. 4. We can see that just by adding the Tele-camera we are already close to the performance of all three. Notably, this camera was occluded by the red filter (Fig. 2) and most examples from qualitative comparison (Fig. 7) showcase how our setup handles discrepancy in red spectral range. In terms of NSE-metric, the combination of Main and Tele even outperforms everything, but statistically insignificantly.
6.4 Effect of Noise in Hyperspectral Camera
Hyperspectral image registration is a process of registering photons, which means our GTs are inherently noisy just like consumer CMOS sensors. Since noise is unpredictable, HSR quality has a physically-reasoned insurmountable boundary. To quantify this boundary, we captured series of HSIs of static scenes with color charts under two different light conditions. The metrics were calculated between a fixed image from a series and the average of remain images, see Tab. 5. The estimated values should be interpreted as approximate bound for any MI-MSR network trained and tested on Doomer dataset. Comparable metric values may indicate overfitting or sensitivity to sensor noise.
@@ 6.4 Effect of Noise in Hyperspectral Camera
Bright scenario (outdoor sunny) 41.32 ± 3.04 0.71 ± 0.11 3.13 ± 0.55
Medium brightness (indoor LED) 37.81 ± 3.49 1.02 ± 0.12 4.46 ± 1.27
Table 5: Metrics measured between several Specim IQ shots of the same scene. These values suggest rough estimates of the highest probable quality of HSR on Doomer.
7 Limitations
The trained models in this work are tightly coupled to the spectral characteristics and quantum efficiency of the specific RGB sensors and optical filters used in our acquisition setup. As a result, deploying the system with different hardware configurations would require collecting a new dataset and retraining the model. This limits immediate out-of-the-box generalization. Future research could explore strategies such as sensor-specific domain adaptation, transfer learning, or SSF-invariant reconstruction frameworks to improve reproducibility and portability across different camera systems.

### v4

@@ 5 Method
The task of MI-HSR involves predicting a hyperspectral image ${\bm{\mathsfit{Y}}}\in\mathbb{R}^{h\times w\times n}$ aligned to a reference RGB image ${\bm{\mathsfit{X}}}_{0}\in\mathbb{R}^{h\times w\times 3}$ given also auxiliary RGB images $\{{\bm{\mathsfit{X}}}_{i}\}_{i=1}^{k-1}\subset\mathbb{R}^{h\times w\times 3}$ from multiple spatially offset sensors. This setup introduces two key challenges: (i) the available hyperspectral supervision $\bm{\mathsfit{Y}}$} is not aligned with any RGB input, and (ii) the input views are misaligned due to differing camera geometries. Our approach (Fig. 4) addresses both issues, one through pre-processing, the other through architectural design.
Figure 4: (Left) Our proposed framework with the Deformable Convolution Alignment Module (DCAM) for solving MI-HSR; any existing HSR network can be plugged. (Right) The proposed GMST++ model featuring GDFN module [45].
5.1 Supervision Warping via Optical Flow
@@
To leverage target hyperspectral image for training, we align it to the reference RGB view using pre-trained optical flow (OF). Since spectral and RGB images differ in modality, we first compute a color projection:
${\bm{C}}_{\mathrm{Y}\to 0}:=\operatorname*{arg\,min}_{{\bm{C}}\in\mathbb{R}^{n\times 3}}\|{\bm{\mathsfit{Y}}}{\bm{C}}-{\bm{\mathsfit{X}}}_{0}\|_{2}^{2}+\mu\|{\bm{C}}\|_{2}^{2}.$
${\bm{\mathsfit{Y}}}{\bm{C}}_{\mathrm{Y}\to 0}$ serves as an approximation of $\bm{\mathsfit{Y}}$} in the color space of the reference RGB. We noted that the approximation becomes noisy in the absence of the regularization term due to ill-posedness of the problem. Since now both views share the same color space, we can estimate a dense correspondence field ${\bm{\mathsfit{D}}}_{0\to\mathrm{Y}}\in\mathbb{R}^{h\times w\times 2}$ using a pre-trained OF model [32]:
${\bm{\mathsfit{D}}}_{0\to\mathrm{Y}}:=\mathrm{OF}({\bm{\mathsfit{X}}}_{0},{\bm{\mathsfit{Y}}}{\bm{C}}_{\mathrm{Y}\to 0}).$
@@
This flow is used to warp the hyperspectral ground truth (GT) to the reference view:
${\bm{\mathsfit{Y}}}_{\mathrm{w}}:=\mathrm{Warp}({\bm{\mathsfit{Y}}},{\bm{\mathsfit{D}}}_{0\to\mathrm{Y}}),\quad{\bm{M}}:=\lfloor\mathrm{Warp}({\bm{E}},{\bm{\mathsfit{D}}}_{0\to\mathrm{Y}})\rfloor,$
@@
where $\bm{M}$} is a binary mask indicating valid visible pixels, $\bm{E}$} is matrix of ones. This enables aligned supervision for training and point-wise loss computation $\mathcal{L}_{1}(\hat{{\bm{\mathsfit{Y}}}}\odot{\bm{M}},{\bm{\mathsfit{Y}}}_{\mathrm{w}}\odot{\bm{M}})$
5.2 Deformable Convolution Alignment
Even with warped supervision, the three input views remain spatially misaligned. Direct flow-based registration is possible but leads to unnecessary error accumulation. Instead, one would prefer leveraging the misaligned views “as-is” in some flow-aware manner. To this end, we make use of deformable convolutions [10], which we feed with dense correspondence field between inputs computed by OF. This allows the convolution layer to deterministically shift each of its 3×3 sampling positions toward geometrically matched locations at an auxiliary view and produce a feature map aligned to the reference view. The kernel weights remain learnable, which allows the filter to handle residual local misalignments and simultaneously learn features useful for further spectral reconstruction.
Similar to the previous section, we first unify color spaces of the reference and auxiliary inputs and compute optical flow between them:
${\bm{C}}_{0\to i}$ $:=\operatorname*{arg\,min}_{{\bm{C}}\in\mathbb{R}^{3\times 3}}\|{\bm{\mathsfit{X}}}_{0}{\bm{C}}-{\bm{\mathsfit{X}}}_{i}\|_{2}^{2},\quad$ $i=1,...,k-1,$
${\bm{\mathsfit{D}}}_{0\to i}$ $:=\mathrm{OF}({\bm{\mathsfit{X}}}_{0}{\bm{C}}_{0\to i},{\bm{\mathsfit{X}}}_{i}),\quad$ $i=1,...,k-1.$
This time we do not need regularization because channels in ${\bm{\mathsfit{X}}}_{0}$ are not as highly correlated as in $\bm{\mathsfit{Y}}$}. The reference input ${\bm{\mathsfit{X}}}_{0}$ is then passed through a standard 3×3 convolution layer, while the auxiliary inputs ${\bm{\mathsfit{X}}}_{i}$ are passed through the aforementioned deformable convolution together with their corresponding deformation fields:
${\bm{\mathsfit{F}}}_{0}:=\mathrm{Conv}({\bm{\mathsfit{X}}}_{0}),\quad{\bm{\mathsfit{F}}}_{i}:=\mathrm{DeformConv}({\bm{\mathsfit{X}}}_{i},{\bm{\mathsfit{D}}}_{0\to i}).$
Because $\mathrm{DeformConv}$ employs a 3×3 kernel, it needs 3×3 sampling points per output pixel. The deformation field supplies only the central offset; the other eight offsets are fixed relative shifts (-1,-1), (-1,0), …, (+1,+1) applied around it. Concatenation of all feature maps $\{{\bm{\mathsfit{F}}}_{i}\}_{i=0}^{k-1}$ serves as an input for Spectral Reconstruction module.
5.3 Spectral Reconstruction
We employ a UNet-like architecture Gated Multi-Stage Transformer (GMST++) heavily inspired by MST++ [5]. Its main building block is Spectral Transformer (ST) block, which in turn consists of Spectral Multi-head Self-attention (S-MSA) and the Gated DConv Feed-forward Network (GDFN) adopted from Restormer [45]. S-MSA treats each 2D slice of the feature tensor as a token and computes attention across channels based on their global correlations. GDFN comprises two branches of successive 1×1 and depth-wise 3×3 convolutions; one branch is GeLU-activated and then multiplied by the other, thereby attenuating some spatial locations and amplifying others. This is particularly helpful in multi-image setting, where small residual misalignments may remain, and their attenuation prevents them from misleading the reconstruction process.
@@ 6 Experiments
Our main intention is to demonstrate the superiority of leveraging a multi-camera imaging system for HSR over reconstruction from a single RGB alone. We evaluate this in both Clean and Real-world settings. The Clean setting consists of simulating perfectly aligned RGB images from a widely recognised HSI benchmark. Such a controlled setup allows us to isolate the effect of increasing the number of cameras and serves as a proof of concept for our proposed acquisition system. The Real-world setting verifies our misalignment-aware framework and shows that the advantages of a multi-camera system can be retained even in an uncontrolled environment.
@@ 6.1 Experimental Setup
Datasets
The Clean setting employs Arad 1K dataset [4] as the source of ground truth HSIs. We simulate three input RGBs given camera response functions of Samsung Galaxy S20, Google Pixel, and iPhone 11 [35]. The spectral filters set was selected specifically for this camera configuration following the methodology described in Sec. 3.2, including leaving the first camera without the filter. Sensor noise was injected according to Gaussian-Poisson model [13]. The Real-world setting employs our Doomer dataset as it is currently the only available dataset suitable for MI-HSR. We split Doomer into train and test in proportion 4:1 and use the authors-provided split for Arad 1K.
Methods
We choose several state-of-the-art HSR methods for comparison with GMST++: AWAN [23], MST++ [5], and SPECAT [43]. Among general image restoration models, we additionally include NAFNet [9]. To evaluate these models in Real-world Multi-image setting, we prepend them with our DCAM module.
Metrics
We adopt the well-known Peak Signal-to-Noise Ratio (PSNR) and Spectral Angle Mapper (SAM). We additionally introduce the Normalized Spectral Error (NSE), which serves as an alternative to MRAE since it does not penalize small values of individual bands and better reflects the integral nature of radiance:
@@ Methods
In the warped GT, some pixels become missing or invalid, as indicated by the mask $\bm{M}$}. Locations where ${\bm{M}}_{i,j}=0$ are excluded from the metric computations.
Implementation details
In all the experiments we used Adam optimizer with learning rate of 0.0004. During training, each image was randomly cropped to 64×64 and optionally flipped vertically or horizontally. Number of epochs was set to 2000 for the Clean setting and to 10000 for the Real-world. Every 30 epochs we ran evaluation loop; the final model is the one with the best MAE on the test set.
@@ 6.2 Results
We trained and tested all HSR methods using two input configurations: single-image $\{{\bm{\mathsfit{X}}}_{0}\}$ and multi-image $\{{\bm{\mathsfit{X}}}_{0},{\bm{\mathsfit{X}}}_{1},{\bm{\mathsfit{X}}}_{2}\}$. The quantitative result for the Clean and Real-world settings are listed in Tables 2 and 3, respectively. Note that we report the number of trainable parameters which means the pre-trained weights of PWC-Net are not counted. However, we do count PWC-Net’s contrubution in GMACs.
Method PSNR, dB ↑ SAM, ° ↓ NSE, % ↓ Params, K GMACs
Single-image input
AWAN [23] 33.91 4.19 7.31 525 25.82
MST++ [5] 33.98 4.06 7.12 546 31.20
SPECAT [43] 33.63 4.17 7.34 346 53.32
NAFNet [9] 33.84 4.14 7.36 372 1.50
GMST++ (ours) 33.99 4.15 7.25 546 31.29
Multi-image input
AWAN [23] 41.98 2.69 4.76 527 44.05
MST++ [5] 41.78 2.56 4.63 547 31.62
SPECAT [43] 39.44 3.05 5.54 350 54.22
NAFNet [9] 41.27 2.63 4.83 373 19.68
GMST++ (ours) 42.17 2.47 4.51 547 31.71
Table 2: Results in Clean setting (Arad 1K dataset). Best and second-best results are highlighted.
Method PSNR, dB ↑ SAM, ° ↓ NSE, % ↓ Params, K GMACs
Single-image input
AWAN [23] 28.65 5.97 12.58 525 25.82
MST++ [5] 28.91 5.23 11.57 655 7.35
SPECAT [43] 27.91 6.09 12.12 389 11.21
NAFNet [9] 29.02 5.39 11.72 372 1.50
GMST++ 28.86 5.28 12.14 654 7.37
Multi-image input
DCAM + AWAN [23] 29.66 4.37 10.45 527 44.06
DCAM + MST++ [5] 30.71 4.19 9.44 657 25.59
DCAM + SPECAT [43] 30.34 4.50 9.72 393 29.54
DCAM + NAFNet [9] 30.53 4.33 9.85 373 19.69
DCAM + GMST++ 31.46 3.91 8.35 658 25.61
Table 3: Results in Real-world setting (Doomer dataset). All methods are supervised by warped GT (${\bm{\mathsfit{Y}}}_{\mathrm{w}}$). To let the methods handle multiple misaligned inputs, we prepend them with our DCAM module. Best and second-best results are highlighted.
Table 2 shows superiority of our GMST++ in multi-image configuration, though it outperforms its nearest competitor only by 0.2 dB. Overall, multi-image input allows substantially better reconstruction than from single-image suggesting that our filtering device design unveils lots of spectral information previously hidden in single RGB inputs.
@@
Figure 5: Qualitative comparison of GMST++ and DCAM + GMST++ predictions under single-image and multi-image input settings respectively. For each scene, a specific spectral band, region of interest, and point of interest are selected. Grayscale renderings correspond to a specific spectral band (indicated in blue on spectral plot). Colorful GT renderings are done with CIE RGB CMFs.
In Real-world setting of Table 3 the MI-HSR by our DCAM + GMST++ method reaches 31.46 PSNR, which is +0.7 dB higher than its nearest competitor and at least +2.44 dB better than any method operating on single-image inputs. The latter improvement secures applicability of multi-image approach in practice. Notably, within the single-image setting our method is far from the best performer, which highlights its tailoring specifically for MI-HSR. Indeed, our motivation for incorporating the GDFN block was to attenuate misaligned regions. In the absence of such misalignments, the model may suppress locally useful information.
We show qualitative performance of our framework based on patch-level comparisons at a selected wavelength as well as radiance profiles at specific points of interest in Fig. 5. The multi-camera system consistently recovers more plausible surface details than the single-camera setup. In the first and third rows, single-camera pipeline misestimates a specific region at a specific wavelength band (730 and 700 nm respectively). In the second row, single-camera pipeline lags behind in overall estimation of brightness, though correctly recovering the shape of the spectrum. These results highlight the improved both spatial and spectral fidelity enabled by multi-view fusion in the MI-HSR system.
@@ 6.3 Ablation Study
To assess importance of each component in our framework, we perform an ablation study by either replacing GDFN with Feed-Forward Network (fallback to MST++) or removing the DCAM module along with the optical flow network in it. All evaluations are performed under Real-world Multi-input setting and are shown in Table 4.
DCAM GDFN PSNR, dB ↑ SAM, ° ↓ NSE, % ↓
28.91 5.23 11.57
+ 30.76 4.41 9.62
+ 30.71 4.19 9.44
+ + 31.46 3.91 8.35
Table 4: Impact of various architecture components on the overall reconstruction quality. Evaluated on the Doomer dataset. Best results are highlighted.
6.4 What is the best possible quality?
Even professional hyperspectral equipment has its own precision, which limits best achievable quality as a GT. Hyperspectral image registration is a process of registering photons, which means our GTs are inherently noisy just like consumer CMOS sensors. To quantify this boundary, we captured series of HSIs of static scenes with color charts under two different light conditions. The metrics were calculated between a fixed image from a series and the average of remain images, see Tab. 5. The estimated values should be interpreted as approximate bound for any MI-MSR network trained and tested on Doomer dataset. Comparable metric values may indicate overfitting or sensitivity to sensor noise.
@@ 6.4 What is the best possible quality?
Bright scenario (outdoor sunny) 41.60 ± 3.41 0.71 ± 0.11 3.00 ± 0.71
Medium brightness (indoor LED) 37.82 ± 3.51 1.02 ± 0.13 4.45 ± 1.28
Table 5: Mean and variance of Specim IQ shots of the same scene.